OpenStudy (anonymous):

ILL GVE MEDAL :3 What is the next number in the following pattern? (Hint: Each number depends on the four numbers that come before it.) 1, 1, 1, 1, 4, 7, 13, 25, 49, 94, , … A.169 B.181 C.196 D.155

Mehek (mehek14):

1+1+1+1=4 1+1+1+4=7 just add the 4 numbers that come before the one you're looking for so \(13+25+49+94=?\)

OpenStudy (anonymous):

181

Mehek (mehek14):

so that's your answer \(\ddot\smile\)

OpenStudy (anonymous):

thx

Mehek (mehek14):

yw

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

one more

Mehek (mehek14):

ok

OpenStudy (anonymous):

14, 26, 38, 50, , …

OpenStudy (anonymous):

A. 58 B. 62 C. 64 D. 56

Mehek (mehek14):

same pattern?

OpenStudy (anonymous):

its a new one

Mehek (mehek14):

any hint?

OpenStudy (anonymous):

c

Mehek (mehek14):

i realized the pattern it's adding 12 each time so 50 + 12 = 62

OpenStudy (anonymous):

thx

Mehek (mehek14):

yw

OpenStudy (anonymous):

your a big help

Mehek (mehek14):

\(\ddot\smile\)